OpenStudy (anonymous):

t=6-3n pls help this one

OpenStudy (anonymous):

with what?

OpenStudy (anonymous):

t=6-3n need to solve this

OpenStudy (anonymous):

you cannot solve an equation having 2 unknowns unless you're given a second equation to accompany it.

OpenStudy (anonymous):

the teacher gave it to me like this i don't know how to do this???

OpenStudy (anonymous):

he or she probably told you to make n the subject of the formula, not to solve the equation.

OpenStudy (anonymous):

do u know how

OpenStudy (anonymous):

yes. \begin{eqnarray*} t&=&6-3n \\ 3n&=&6-t \\ n&=&\frac{6-t}{3} = 2-\frac{t}{3} \end{eqnarray*}

OpenStudy (anonymous):

what about this one t=12n

OpenStudy (anonymous):

thx for the other one

OpenStudy (anonymous):

just divide by 12: \[\frac{t}{12}=n\]

OpenStudy (anonymous):

now my head is more clear

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

what about this t=6-3square n

OpenStudy (anonymous):

you mean \[t=6-3n^2\]?

OpenStudy (anonymous):

is on the 3

OpenStudy (anonymous):

\[t=6-3^2n\] like this?

OpenStudy (anonymous):

yes, how do u do that 2on this

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\begin{eqnarray*} 3^2n&=&6-t \\ 9n&=&6-t \\ n&=&\frac{6-t}{9} \end{eqnarray*}

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

t=2n-2